Titanium is a transition metal used in many alloys because it is extremely strong and lightweight. Titanium tetrachloride (TiCl4) is extracted from titanium oxide (TiO2) using chlorine and carbon:TiO2(s) + C(s) + Cl2(g) → TiCl4(s) + CO2(g)1.1Calculate the mass of Cl2 gas needed to react with 1.25 moles of TiO2.(4)1.2Say 2.45 grams of TiO2 reacted with 3.25 grams of chlorine gas. During this reaction 4.35 grams of TiCl4was produced. Identify and quantify the limiting reagent. (5)1.3Calculate the amount of excess reagent remaining on completion of the reaction


1
Expert's answer
2021-12-13T17:09:03-0500

The reaction is represented by the following equation:

TiO2 + 2Cl2 + C --> TiCl4 + CO2

Therefore,

"m(Cl_2)=1.25mol(TiO_2)\times\frac{2mol(Cl_2)}{1mol(TiO_2)}\times\frac{70.91g(Cl_2)}{1mol(Cl_2)}=177g"


Answer: 177 g